# Pulse-train propagation in nonlinear Kerr media governed by higher-order dispersion

### v1

@@ IV Conclusion
In conclusion, we have discovered three classes of pulse-trains whose amplitudes, widths and wavenumbers are different with the distinctive property of exhibiting a relation between the amplitude and duration determined by the sign of a joint parameter uniquely. We have found that these structures have the same velocity, which depends only on the three dispersion parameters. It is demonstrated that higher-order dispersion may lead to the formation of double-humped pulse-trains besides the ones taking the usual form with single peak in each periodic unit. It is also shown that the pulse-trains are obtained without necessarily assuming a specified condition on the fiber parameters. The results presented here may be useful for experimental realization of pulse-trains in optical fibers and may help in stimulating more research in understanding their optical transmission properties.

### v2

@@ VI Conclusion
In conclusion, we have discovered three classes of pulse-trains whose amplitudes, widths and wavenumbers are different with the distinctive property of exhibiting a relation between the amplitude and duration determined by the sign of a joint parameter uniquely. We have found that these structures have the same velocity, which depends only on the three dispersion parameters. It is demonstrated that higher-order dispersion may lead to the formation of double-humped pulse-trains besides the ones taking the usual form with single peak in each periodic unit. It is also shown that the pulse-trains are obtained without necessarily assuming a specified condition on the fiber parameters. Due to the precise nature, the nonlinear waves presented here may be profitably used in designing the optimal fiber system experiments. These results can be exploited to experimentally realize pulse-trains in fiber systems and may help in stimulating more research in understanding their optical transmission properties.